OpenStudy (anonymous):

Which fraction is between 1/2 and 3/4 on a number line? A. 3/8 B. 2/5 C. 3/5 D. 7/8

OpenStudy (anonymous):

c use your calculator to check

OpenStudy (anonymous):

\[\left\{\frac{1}{2},\frac{3}{4}\right\}=\left\{\frac{4}{8},\frac{6}{8}\right\} \]Looks like 5/8 to me.

OpenStudy (anonymous):

It's not one of the choices, but you're right.

OpenStudy (anonymous):

its 3/5,change all to fractions with 40 as denominator

OpenStudy (anonymous):

\[\left\{\frac{1}{2},\frac{3}{4}\right\}=\left\{\frac{20}{40},\frac{30}{40}\right\} \]Looks like 25/40 or 5/8 to me.